# Batch Predictive Inference

### v4

@@ 1.3 Related work
The idea of distribution-free prediction sets dates back at least to the pioneering works of Wilks (1941), Wald (1943), Scheffe and Tukey (1945), and Tukey (1947, 1948). Distribution-free inference has been extensively studied in recent works (see, e.g., Saunders et al., 1999; Vovk et al., 1999; Papadopoulos et al., 2002; Vovk et al., 2005; Vovk, 2013a; Lei et al., 2013; Lei and Wasserman, 2014; Lei et al., 2018; Angelopoulos et al., 2023; Guan, 2023; Romano et al., 2020; Liang et al., 2023; Dobriban and Yu, 2023). Predictive inference methods (e.g., Geisser, 2017, etc) have been developed under various assumptions (see, e.g., Bates et al., 2021; Park et al., 2022a, b; Sesia et al., 2023; Qiu et al., 2023; Li et al., 2022; Kaur et al., 2022; Si et al., 2024; Lee et al., 2024). Overviews of the field are provided by Vovk et al. (2005); Shafer and Vovk (2008), and Angelopoulos et al. (2023). For exchangeable data, conformal prediction and split conformal prediction (Vovk et al., 2005; Papadopoulos et al., 2002) provide a general framework for distribution-free predictive inference.
Distribution-free predictive inference for multiple test points has been extensively studied in the context of outlier detection and selection (Bates et al., 2023; Jin and Candès, 2023b, a; Gui et al., 2024). These works apply multiple testing methods to conformal p-values for inference on multiple test outcomes. Vovk (2013c) discuss transductive conformal methods for constructing a prediction region for the vector of test outcomes. Lee et al. (2024) introduces a method for constructing simultaneous prediction sets for multiple outcomes under covariate shift with a conditional guarantee.
Gazin et al. (2024) study a closely related problem setting to our paper, transductive conformal inference with adaptive scores. In this scenario, they derive the joint distribution of multiple test conformal p-values in the case of no ties between non-conformity scores, which is equivalent to the joint distribution of their ranks, and which we use in the proof of our Theorem 1. Gazin et al. (2024) also give intriguing equivalent characterizations of this distribution, for instance in terms of Pólya urns (see also Gazin (2024) for a functional CLT for the coverage). Further, they apply these results to several problems, including controlling the false coverage rate of the prediction sets for multiple test points. For this probleem Marques F. (2025) derived the distribution of the coverage. This problem is also considered one of our use cases in this work, and we will provide further discussion in Section 3.1.
@@ 2 The batch PI methodology
Here and below, we will suppose that the calibration and test data $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n}),(X_{n+1},Y_{n+1}),\ldots,(X_{n+m},Y_{n+m})$ are exchangeable, unless explicitly specified otherwise. If $m=1$, i.e., we have only one test point, then the coverage guarantee (1) can be achieved simply by standard distribution-free prediction methods, such as full and split conformal prediction (Vovk et al., 2005; Papadopoulos et al., 2002), for any function $g$. For example, if we set $g(z)$ as the nonconformity score, i.e., $g({z})=|y-\hat{\mu}(x)|$, for all $z=(x,y)$, then the condition (1) is equivalent to $\mathbb{P}\left\{{s(X_{n+1},Y_{n+1})\in\widehat{C}(\mathcal{D}_{n})}\right\}\geqslant 1-\alpha$, and split conformal prediction attains this guarantee with $\widehat{C}(\mathcal{D}_{n})=\Big{(}-\infty,Q_{1-\alpha}\left(\sum_{i=1}^{n}\frac{1}{n+1}\delta_{s(X_{i},Y_{i})}+\frac{1}{n+1}\delta_{\infty}\right)\Big{]}$ (Saunders et al., 1999; Vovk et al., 1999, 2005; Papadopoulos et al., 2002).
However, for multiple test points, it turns out that constructing a useful distribution-free prediction set that satisfies (1) is a nontrivial task. One can imagine a number of direct approaches (See Appendix A for details). For instance, one could partition the calibration dataset into non-overlapping subsets of the same size as the test data, and then use standard conformal prediction for the function $g$ applied to the partitions and the test data. However, this approach produces nontrivial or short prediction sets only when the calibration set size is much larger than the test set size. For example, if $n<m(1/\alpha-1)$, it leads to a trivial prediction set (See Section A.1). A direct extension of split conformal prediction has similar issues (See Section A.2). Finally, a direct extension of full conformal prediction is computationally prohibitive (See Section A.3). Thus, none of these direct approaches are practically viable in the setting we are interested in and therefore will not be given further consideration.
@@ 2.1 Proposed method: batch PI
In this section, we introduce our new batch PI procedure, which can be less conservative and more computationally efficient than the naive methods described above. To introduce our method, it is helpful to review the idea of split conformal prediction. Suppose we have only one input $X_{n+1}$ beyond the calibration data $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})$. The first step is to construct a nonconformity score $s:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}_{\geqslant 0}$; based on data that is independent of the calibration and test datasets. Let us write $S_{i}=s(X_{i},Y_{i})$ for $i\in[n+1]$. The split conformal prediction set is given by
@@
It is known that if $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n}),(X_{n+1},Y_{n+1})$ are exchangeable, the prediction set $\widehat{C}_{n}$ from (4) satisfies the following coverage guarantee (Vovk et al., 2005): $\mathbb{P}\left\{{Y_{n+1}\in\widehat{C}_{n}(X_{n+1})}\right\}\geqslant 1-\alpha$.
The key intuition is as follows: Let $S_{(1)},S_{(2)},\ldots,S_{(n)}$ be the order statistics of $S_{1},S_{2},\ldots,S_{n}$, breaking ties uniformly at random. Then, the rank $R\in\{1,\ldots,n+1\}$ such that $S_{(R)}$ is the smallest upper bound among the observed scores for the unobserved score $S_{n+1}$ follows a uniform distribution over $\{1,\ldots,n+1\}$. (where we define $S_{(n+1)}=+\infty$) Then, because $Y_{n+1}\in\widehat{C}_{n}(X_{n+1})$ is implied by $R\leqslant\lceil(1-\alpha)(n+1)\rceil$, the coverage probability is at least $1-\alpha$.
The batch PI method.
@@
Note that $|H|=\binom{n+m}{m}$. This set will represent the ranks of the test data points among the calibration data points.
Moreover, we also need a way to order these ranks. In the standard conformal case where $m=1$, the ranks are ordered as $11\leqslant\ldots\leqslant n+$ 1, but for our case there is no default ordering. Hence to allow for the maximum flexibility, we introduce a general rank-ordering function $\tilde{h}:H\rightarrow\mathbb{R}$ that we will use to prioritize the ranks. We will later discuss at length the choice of this function.
@@
$q_{L}=Q_{\beta}^{\prime}\left(\sum_{r_{1:m}\in H}\frac{1}{\binom{n+m}{m}}\delta_{\tilde{h}(r_{1:m})}\right),\quad q_{U}=Q_{1-\gamma}\left(\sum_{r_{1:m}\in H}\frac{1}{\binom{n+m}{m}}\delta_{\tilde{h}(r_{1:m})}\right).$ (6)
By definition, if $R_{1:m}$ is distributed uniformly over $H$, then $\mathbb{P}\left\{{\tilde{h}(R_{1:m})\in[q_{L},q_{U}]}\right\}\geqslant 1-\alpha$. However, since we are interested in covering the values of the function $h$ (or equivalently $g$), we also need a way to define an appropriate range of $h$ values. We do this by first considering the pre-image of $[q_{L},q_{U}]$ under $\tilde{h}$, and then considering its image under $h$. It turns out that we also need to consider certain corner cases (e.g., when the rank is $n+1$), and so with $S_{(0)}=\inf s$ and $S_{(n+1)}=\sup s$, we define
@@
See Algorithm 1. For completeness, we also provide a one-sided version of the batch PI prediction set algorithm, which simplifies slightly, see Algorithm 8 in Section B.1. The validity of batch PI is proved in Theorem 1.
@@
The proof is deferred to Appendix E, and here we offer some intuition. Suppose the scores $S_{1},\ldots,S_{n+m}$ are distinct almost surely, and define $R_{n+1},R_{n+2},\ldots,R_{n+m}$ as $R_{n+j}=\min\{r\in\{1,2,\ldots,n\}:S_{(r)}\geqslant S_{n+j}\},\textnormal{ for}j=1,2,\ldots,m$, where we let $R_{n+j}=n+1$ if $S_{(n)}<S_{n+j}$. Let $R_{(n+1)},\ldots,R_{(n+m)}$ be their order statistics. Through the exchangeability condition, it follows that $(R_{(n+1)},\ldots,R_{(n+m)})\sim\textnormal{Unif}(H)$ (Gazin et al., 2024). This implies that for any fixed subset $I$ of $H$ with $|I|\geqslant(1-\alpha)|H|,$ we have $\mathbb{P}\left\{{(R_{(n+1)},\ldots,R_{(n+m)})\in I}\right\}\geqslant 1-\alpha$. By construction, the set $\{r_{1:m}\in H,\tilde{h}(r_{1:m})\in[q_{L},q_{U}]\}$ satisfies this condition, and thus we have
$\mathbb{P}\left\{{h(S_{(n+1)},\ldots,S_{(n+m)})\leqslant[B_{L},B_{U}]}\right\}\geqslant\mathbb{P}\left\{{h(S_{R_{(n+1)}},\ldots,S_{R_{(n+m)}})\leqslant[B_{L},B_{U}]}\right\}\geqslant 1-\gamma,$
@@
Since $\binom{n+m}{m}\sim(1+n/m)^{m}$, the computational cost of a direct enumeration-based approach for batch PI can be extremely high when the number of the calibration and test data points are large.
To confirm that this computation is indeed hard in general, we take the perspective of standard computational complexity theory (e.g., Garey and Johnson, 1979), where the difficulty of problems is characterized according to the number of steps it takes to execute them on a standard model of computation called the Turing machine. Tractable problems usually have a polynomial running time, while there is a potentially broader class of problems—called NP—whose solutions can be verified in polynomial time. There is a large set of difficult combinatorial problems— called NP-hard problem—that are at least as hard as any problem in NP. By showing that solving the prediction set problem can be used to solve the so-called vertex cover problem (e.g., Garey and Johnson, 1979), we show that computing batch PI is in general NP-hard.
@@
For the choice of the rank ordering function $\tilde{h}$, we have the following considerations. To ensure validity, this function cannot depend on the calibration scores $S_{1},\ldots,S_{n}$. However, to obtain a short and informative prediction sets, the values $h(S_{(r_{1})},\ldots,S_{(r_{m})})$ when varying $(r_{1},\ldots,r_{m})$ should be similarly ordered as the values $\tilde{h}(r_{1},\ldots,r_{m})$. We will describe below two heuristic strategies to achieve this goal.

### v5

@@ 1.3 Related work
The idea of distribution-free prediction sets dates back at least to the pioneering works of Wilks (1941), Wald (1943), Scheffe and Tukey (1945), and Tukey (1947, 1948). Distribution-free inference has been extensively studied in recent works (see, e.g., Saunders et al., 1999; Vovk et al., 1999; Papadopoulos et al., 2002; Vovk et al., 2005; Vovk, 2013a; Lei et al., 2013; Lei and Wasserman, 2014; Lei et al., 2018; Angelopoulos et al., 2023; Guan, 2023; Romano et al., 2020; Liang et al., 2023; Dobriban and Yu, 2025). Predictive inference methods (e.g., Geisser, 2017, etc) have been developed under various assumptions (see, e.g., Bates et al., 2021; Park et al., 2022a, b; Sesia et al., 2023; Qiu et al., 2023; Li et al., 2022; Kaur et al., 2022; Si et al., 2024; Lee et al., 2024). Overviews of the field are provided by Vovk et al. (2005); Shafer and Vovk (2008), and Angelopoulos et al. (2023). For exchangeable data, conformal prediction and split conformal prediction (Vovk et al., 2005; Papadopoulos et al., 2002) provide a general framework for distribution-free predictive inference.
Distribution-free predictive inference for multiple test points has been extensively studied in the context of outlier detection and selection (Bates et al., 2023; Jin and Candès, 2023b, a; Gui et al., 2024). These works apply multiple testing methods to conformal p-values for inference on multiple test outcomes. Vovk (2013c) discuss transductive conformal methods for constructing a prediction region for the vector of test outcomes, where transductive means that the predictor (inducing the non-conformity score) used to construct the prediction sets can depend on the test dataset. Lee et al. (2024) introduces a method for constructing simultaneous prediction sets for multiple outcomes under covariate shift with a conditional guarantee.
Gazin et al. (2024) study a closely related problem setting to our paper, transductive conformal inference with adaptive scores. In this scenario, they derive the joint distribution of multiple test conformal p-values in the case of no ties between non-conformity scores, which is equivalent to the joint distribution of their ranks, and which we use in the proof of our Theorem 1. Gazin et al. (2024) also give intriguing equivalent characterizations of this distribution, for instance in terms of Pólya urns (see also Gazin (2024) for a functional CLT for the coverage). Further, they apply these results to several problems, including controlling the false coverage rate of the prediction sets for multiple test points. For this problem, Marques F. (2025) derived the distribution of the coverage. This problem is also considered in one of our use cases in this work, and we will provide further discussion in Section 3.2.
@@ 2 Main results
Here and below, we will suppose that the calibration and test data $(X_{1},Y_{1}),\ldots,(X_{n+m},Y_{n+m})$ are exchangeable, unless explicitly specified otherwise. If $m=1$, i.e., we have only one test point, then the coverage guarantee (1) can be achieved simply by standard distribution-free prediction methods, such as full and split conformal prediction (Vovk et al., 2005; Papadopoulos et al., 2002), for any function $g$. For example, if we set $g(z)$ as the nonconformity score, i.e., $g({z})=|y-\hat{\mu}(x)|$, for all $z=(x,y)$, then the condition (1) is equivalent to the standard marginal coverage guarantee $\mathbb{P}\big{\{}s(X_{n+1},Y_{n+1})\in\widehat{C}(\mathcal{D}_{n})\big{\}}\geqslant 1-\alpha$, and split conformal prediction attains this guarantee with the following prediction set (Saunders et al., 1999; Vovk et al., 1999, 2005; Papadopoulos et al., 2002).
$\widehat{C}(\mathcal{D}_{n})=\Bigg{(}-\infty,Q_{1-\alpha}\left(\sum_{i=1}^{n}\frac{1}{n+1}\delta_{s(X_{i},Y_{i})}+\frac{1}{n+1}\delta_{\infty}\right)\Bigg{]}.$
However, for multiple test points, it turns out that constructing a useful distribution-free prediction set that satisfies (1) is a nontrivial task. One can imagine a number of direct approaches, such as directly extending split conformal or full conformal prediction; however, it turns out that their usefulness is limited to a small range of settings, as we discuss next. The reader may directly skip to Section 2.2 to read the description of our proposed method.
2.1 Baseline approaches
In this Section, we consider several possible reasonable alternative approaches to our Batch PI approach introduced in the next Section, and we discuss their limitations.
2.1.1 Partitioning the calibration data
A potential approach to achieve (1) is to partition the calibration data, to obtain multiple groups of observations that are exchangeable with the test set. Specifically, suppose $n=mq+r$ where $q$ is a non-negative integer and $10\leqslant r\leqslant m-$ 1. Let
$\tilde{Z}_{k}=\{Z_{(k-1)m+1},Z_{(k-1)m+2},\ldots,Z_{km}\}\text{ for $k\in[q]$ and }\tilde{Z}_{\textnormal{test}}=\{Z_{n+1},\ldots,Z_{n+m}\}.$
Then it is clear that $g(\tilde{Z}_{1}),\ldots,g(\tilde{Z}_{q}),g(\tilde{Z}_{\textnormal{test}})$ are exchangeable, and thus we can apply split conformal prediction to obtain the following prediction set for $g(\tilde{Z}_{\textnormal{test}})$:
$\widehat{C}(\mathcal{D}_{n})=\left[Q_{\beta}^{\prime}\left(\sum_{k=1}^{q}\frac{1}{q+1}\delta_{g(\tilde{Z}_{k})}+\frac{1}{q+1}\delta_{-\infty}\right),Q_{1-\gamma}\left(\sum_{k=1}^{q}\frac{1}{q+1}\delta_{g(\tilde{Z}_{k})}+\frac{1}{q+1}\delta_{\infty}\right)\right],$
2.1.2 Extending split conformal prediction
where $S_{i}=s(Z_{i})$. For $s_{1}\leqslant s_{2}\leqslant\ldots\leqslant s_{m}$, we define $h(s_{1},s_{2},\ldots,s_{m})$ as $\sup h$ if $s_{m}=\sup s$ and $h$ is not well-defined, e.g., $\sup s=+\infty$ and $h(s_{1},\ldots,s_{m})=\sum_{j=1}^{m}s_{i}$. Similarly, we define $h(s_{1},s_{2},\ldots,s_{m})$ as $\inf h$ if $s_{1}=\inf s$ and $h$ is not well-defined; while noting that only one of the two cases can occur below. Then, the adapted split conformal prediction set $\widehat{C}(\mathcal{D}_{n})$ is defined as:
However, this approach still faces limitations unless $n\gg m$. For instance, consider the scenario where $n=m$ Then half of the $(\bar{S}_{i})_{1\leqslant i\leqslant n+m}$ values are $\sup s$, likely leading to a near-trivial upper bound in (6).
2.1.3 Split conformal prediction with Bonferroni correction
Alternatively, one may consider bounding individual scores and then combining them using a Bonferroni-type approach. Specifically, let $\hat{q}_{\beta/m}^{\prime}$ and $\hat{q}_{1-\gamma/m}$ denote the lower and upper score bounds obtained from split conformal prediction, using the adjusted level $\beta/m$ and $\gamma/m$ (where $\beta+\gamma=\alpha$):
$\hat{q}_{\beta/m}^{\prime}=Q_{\beta/m}^{\prime}\left(\sum_{i=1}^{n}\tfrac{1}{n+1}\delta_{S_{i}}+\tfrac{1}{n+1}\delta_{\inf s}\right),\hat{q}_{1-\gamma/m}=Q_{1-\gamma/m}\left(\sum_{i=1}^{n}\tfrac{1}{n+1}\delta_{S_{i}}+\tfrac{1}{n+1}\delta_{\sup s}\right).$
Then the following prediction set attains the coverage guarantee at level $1-\alpha$:
$\left[h(\hat{q}_{\beta/m}^{\prime},\hat{q}_{\beta/m}^{\prime},\cdots,\hat{q}_{\beta/m}^{\prime}),\;h(\hat{q}_{1-\gamma/m},\hat{q}_{1-\gamma/m},\cdots,\hat{q}_{1-\gamma/m})\right].$
The proof follows directly from the union bound and the monotonicity of $h$. This method suffers from issues similar to the previous ones: unless $n>m/\alpha$ (i.e., $n\gg m$), we have $\hat{q}_{\beta/m}^{\prime}=\inf s$ and $\hat{q}_{1-\gamma/m}=\sup s$, which lead to a trivial prediction set.
2.1.4 Extending full conformal prediction
To avoid the issue of having a large mass at $\infty$ or $\infty$- ∞, one may try to construct a full conformal-type prediction set instead of relying on split conformal-type constructions. For example, we can first construct a joint prediction set $\widehat{C}_{n}(X_{n+1},\ldots,X_{n+m})$ for $(y_{(n+1):(n+m)})$ as
[…]
where $S_{i}^{\tilde{y}}=s^{\tilde{y}}(X_{i},Y_{i})$ and $s^{\tilde{y}}$ is the nonconformity score constructed from $(X_{1},Y_{1}),\ldots,$ $(X_{n},Y_{n})$ and $(X_{n+1},y_{n+1}),\ldots,(X_{n+m},y_{n+m})$—this step is essentially equivalent to the transductive conformal prediction Vovk (2013c). Then the prediction set for $g(\{Z_{n+1},\ldots,Z_{n+m}\})$ can be constructed as
[…]
To summarize, none of these direct approaches are practically viable in the setting we are interested in—in terms of the usefulness of the prediction set or the computational burden—and therefore will not be given further consideration.
@@ 2.2 Proposed method: batch PI
In this section, we introduce our batch PI procedure, which can be less conservative and more computationally efficient than the baseline methods described above. To introduce our method, it is helpful to review the idea of split conformal prediction. Suppose we have only one test input $X_{n+1}$. The first step is to construct a nonconformity score function $s:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$; based on data that is independent of the calibration and test datasets. Let us write $S_{i}=s(X_{i},Y_{i})$ for $i\in[n+1]$. The split conformal prediction set is given by
@@
It is known that if $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n}),(X_{n+1},Y_{n+1})$ are exchangeable, the prediction set $\widehat{C}_{n}$ from (8) satisfies the following coverage guarantee (Vovk et al., 2005): $\mathbb{P}\big{\{}Y_{n+1}\in\widehat{C}_{n}(X_{n+1})\big{\}}\geqslant 1-\alpha.$
The key intuition is as follows: Let $S_{(1)},\ldots,S_{(n)}$ be the order statistics of $S_{1},\ldots,S_{n}$, breaking ties uniformly at random. Then, the rank $R\in[n+1]$ such that $S_{(R)}$ is the smallest upper bound among the observed scores for the unobserved score $S_{n+1}$ follows a uniform distribution over $[n+1]$; where we define $S_{(n+1)}=+\infty$. Then, because $Y_{n+1}\in\widehat{C}_{n}(X_{n+1})$ is implied by $R\leqslant\lceil(1-\alpha)(n+1)\rceil$, the coverage probability is at least $1-\alpha$.
@@
Note that $|H|=_{n+1}\mathrm{H}_{m}=\binom{n+m}{m}$. This set will represent the ranks of the test data points among the calibration data points666 Denoting the order statistics of the test scores $S_{n+1},\cdots,S_{n+m}$ as $S_{(1)}^{\text{test}},\cdots,S_{(m)}^{\text{test}}$, our strategy is to bound each order statistic—which is unobserved—by one of the observed scores. Let $S_{(1)},\cdots S_{(n)}$ be the order statistics of the calibration scores which we have access to. Now, for each $j=1,2,\cdots,m$, we consider the smallest (observed) $S_{(r_{j})}$ which bounds (the unobserved) $S_{(j)}^{\text{test}}$. We can then bound our target as well. For example, if we are interested in the mean, we leverage $({S_{(1)}^{\text{test}}+\cdots+S_{(m)}^{\text{test}}})/{m}\leqslant({S_{(r_{1})}+\cdots+S_{(r_{m})}})/{m}.$ This motivates the definition of $H$ as the set of ranks the test scores..
Moreover, we also need a way to order these ranks. In the standard conformal case where $m=1$, the ranks are ordered as $11\leqslant\ldots\leqslant n+$ 1, but for our case, there is no default ordering. Hence to allow for the maximum flexibility, we introduce a general rank-ordering function $\tilde{h}:H\rightarrow\mathbb{R}$ that we will use to prioritize the ranks. We will later discuss at length the choice of this function.
@@
$q_{L}=Q_{\beta}^{\prime}\bigg{(}\sum_{r_{1:m}\in H}\frac{1}{\binom{n+m}{m}}\delta_{\tilde{h}(r_{1:m})}\bigg{)},\quad q_{U}=Q_{1-\gamma}\bigg{(}\sum_{r_{1:m}\in H}\frac{1}{\binom{n+m}{m}}\delta_{\tilde{h}(r_{1:m})}\bigg{)}.$ (10)
By definition, if $R_{1:m}$ is distributed uniformly over $H$, then $\mathbb{P}\big{\{}\tilde{h}(R_{1:m})\in[q_{L},q_{U}]\big{\}}\geqslant 1-\alpha$. However, since we are interested in covering the values of the function $h$ (or equivalently $g$), we also need a way to define an appropriate range of $h$ values. We do this by first considering the pre-image of $[q_{L},q_{U}]$ under $\tilde{h}$, and then considering its image under $h$. It turns out that we also need to consider certain corner cases (e.g., when the rank is $n+1$), and so with $S_{(0)}=\inf s$ and $S_{(n+1)}=\sup s$ [^7] The notations $S_{(0)}$ and $S_{(n+1)}$ are introduced solely for notational convenience in the expressions for $B_{L}$ and $B_{U}$, and they do not correspond to actual order statistics., we define
@@
See Algorithm 1. For completeness, we also provide a one-sided version of the batch PI prediction set algorithm, which simplifies slightly, see Algorithm 3. The validity of batch PI is proved in Theorem 1.
@@
The proof is deferred to the Appendix, and here we offer some intuition. Suppose the scores $S_{1},\ldots,S_{n+m}$ are distinct almost surely [^8] Almost sure distinctness is not an assumption of the theorem; it is assumed here solely for simplicity in the intuitive proof sketch., and define $R_{n+1},\ldots,R_{n+m}$ as
$R_{n+j}=\min\{r\in\{1,2,\ldots,n\}:S_{(r)}\geqslant S_{n+j}\},\textnormal{ for}j\in[m],$
where we let $R_{n+j}=n+1$ if $S_{(n)}<S_{n+j}$. Let $(R_{(n+j)})_{j\in[m]}$ be their order statistics. Through the exchangeability condition, it follows that $(R_{(n+1)},\ldots,R_{(n+m)})\sim\textnormal{Unif}(H)$. Thus, for any subset $I$ of $H$ with $|I|\geqslant(1-\alpha)|H|,$$\mathbb{P}\left\{{(R_{(n+1)},\ldots,R_{(n+m)})\in I}\right\}$ $\geqslant 1-\alpha$.
Denoting the $j$-th order statistics in $S_{n+1},\cdots,S_{n+m}$ as $S_{(j)}^{\text{test}}$ for $j\in[m]$, we thus have by construction that
$\mathbb{P}\left\{{h(S_{(1)}^{\text{test}},\ldots,S_{(m)}^{\text{test}})\in[B_{L},B_{U}]}\right\}\geqslant\mathbb{P}\left\{{h(S_{R_{(n+1)}},\ldots,S_{R_{(n+m)}})\in[B_{L},B_{U}]}\right\}\geqslant 1-\gamma,$
@@
Since $\binom{n+m}{m}\sim(1+n/m)^{m}$, the computational cost of an enumeration-based approach for batch PI can be extremely high when there are many calibration and test datapoints.
To confirm that this computation is indeed hard in general, we take the perspective of standard computational complexity theory (e.g., Garey and Johnson, 1979), where the difficulty of problems is characterized according to the number of steps it takes to execute them on a standard model of computation called the Turing machine. Tractable problems usually have a polynomial running time, while there is a potentially broader class of problems—called NP—whose solutions can be verified in polynomial time. There is a large set of difficult combinatorial problems—called NP-hard problems—that are at least as hard as any problem in NP. By showing that solving the prediction set problem can be used to solve the so-called vertex cover problem (e.g., Garey and Johnson, 1979), we show that computing batch PI is in general NP-hard.
@@
For the choice of the rank ordering function $\tilde{h}$, we have the following considerations. To ensure validity, this function cannot depend on the calibration scores $S_{1},\ldots,S_{n}$. However, to obtain a short and informative prediction sets, the values $h(S_{(r_{1})},\ldots,S_{(r_{m})})$ when varying $(r_{1},\ldots,r_{m})$ should be similarly ordered as the values $\tilde{h}(r_{1},\ldots,r_{m})$. To see this, observe that the upper bound $B_{U}$ in (11) is, roughly speaking, defined as the “maximum of the $h$ values among those with small $\tilde{h}$ values”. We describe below two heuristic strategies to achieve this goal, and provide a more detailed discussion in Appendix A.